OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x - 3y - 5 = 0 2x + 3y - 5 = 0 -2x + 3y - 5 = 0

OpenStudy (anonymous):

m = -4/6 = -2/3 equation y - 3=-2/3(x + 2) 3y - 9 = -2x-4 2x+3y -5 = 0